# Causal Inference for a Hidden Treatment

### v2

@@ 6 Extention to other functionals
$\begin{split}\phi_{\mathrm{g}}^{F}(Y,A^{*},X;\psi^{*})&=\sum_{a^{*}=0,1}IF_{a^{*};\psi^{*}}=\sum_{a^{*}=0,1}\Bigl{\{}q_{0,a^{*}}(X)h_{0,a^{*}}(X)g_{1,a^{*}}(Y,X)\mathbb{I}(A^{*}=a^{*})\\ &\quad+q_{0,a^{*}}(X)g_{2,a^{*}}(Y,X)\mathbb{I}(A^{*}=a^{*})+h_{0,a^{*}}(X)g_{3,a^{*}}(Y,X)+g_{4,a^{*}}(Y,X)\Bigr{\}},\end{split}$ (21)
where $q_{0,a^{*}}(X)$ and $h_{0,a^{*}}(X)$ are unknown nuisance functions, $g_{j,a^{*}}$ are known functions of $(Y,X)$, and $IF_{a^{*};\psi^{*}},a^{*}=0,1$, belongs to the doubly robust class of influence functions studied in Robins et al. (2008). To simplify the exposition, we suppress the influence function dependence on $\psi^{*}$ throughout as $E[IF_{a^{*};\psi^{*}}]=0,a^{*}=0,1$. The next theorem characterizes for each influence function in the class above, a corresponding observed data influence function in the hidden treatment model. While we state the result for the class given above, the proof is more general as it considers an even larger class of doubly robust influence functions introduced in Ghassami et al. (2022), an important generalization which includes influence functions for causal functionals identified using the proximal causal inference framework (Miao et al., 2016, 2018; Tchetgen Tchetgen et al., 2024) that do not strictly fall within the above doubly robust class (Cui et al., 2023). The proof is deferred to the Supplementary Materials §LABEL:supp-sec:proof_extension.
@@
Suppose that the influence functions for the parameter of interest $\psi$ in the full data model $\mathcal{M}^{*}$ are of the form in (21), indexed by functions $\mathrm{g}=(g_{1,a^{*}},g_{2,a^{*}},g_{3,a^{*}},g_{4,a^{*}},a^{*}=0,1)\in\mathcal{G}$ for a certain space of functions $\mathcal{G}$ in $L_{0}^{2}(X,Y)$ implied by the semiparametric model; then
@@
is an influence function for $\psi$ in the hidden treatment model $\mathcal{M}$, where
$\begin{split}M_{a^{*},A}=\frac{A-E(A\mid A^{*}=1-a^{*},X)}{E(A\mid A^{*}=a^{*},X)-E(A\mid A^{*}=1-a^{*},X)},\\ M_{a^{*},Z}=\frac{Z-E(Z\mid A^{*}=1-a^{*},X)}{E(Z\mid A^{*}=a^{*},X)-E(Z\mid A^{*}=1-a^{*},X)}.\end{split}$ (23)
@@
    Perform a one-step update in the direction of estimated efficient influence function
    $\hat{\psi}^{eff}=\hat{\psi}+\mathbb{P}_{n}\hat{\phi}_{eff}(Y,A,Z,X)$
    where $\hat{\phi}_{eff}(Y,A,Z,X)$ is an estimator of the efficient influence function of ${\phi}_{eff}(Y,A,Z,X)=\Pi(\phi_{\mathrm{g}}\mid\Lambda)$ evaluated at $\hat{\psi}$ given in (LABEL:supp-eq:est_eif) in the Supplementary Materials §LABEL:supp-sec:closed_form_w.
@@
Semi-linear structural regression model (continued). The full data influence function in (5) can be reformulated as
$\phi^{F}_{sl}$ $=\mathbb{I}(A^{*}=1)h(X)\{Y-\beta_{a^{*}}-\beta_{x}(X)\}\{1-\textrm{pr}(A^{*}=1\mid X)\}$
$\quad+\mathbb{I}(A^{*}=0)h(X)\{Y-\beta_{x}(X)\}\{-\textrm{pr}(A^{*}=1\mid X)\}.$
An influence function in the observed data model can be obtained by replacing $\mathbb{I}(A^{*}=a^{*}),a^{*}=0,1$ with $M_{a^{*},A}(A,X)M_{a^{*},Z}(Z,X)$ in (23), which gives
@@ 7 Simulation study
In this section, we evaluate the performance of our proposed estimator for both binary and continuous outcomes. The covariates $X=(X_{1},X_{2})$ are generated from a 2-dimensional uniform $[0,1]^{2}$ distribution. We then generate $A^{*}$ conditional on $X$ from a Bernoulli distribution with $\textrm{pr}(A^{*}=1\mid X)=\text{logit}^{-1}(0.7+0.5X_{1}+0.5X_{2})$ where $\text{logit}(p)=\ln p(1-p)^{-1}$. Next, we generate $A$ and $Z$ from
$\textrm{pr}(A=1\mid A^{*}=1,X)$ $=\text{logit}^{-1}(1+2X_{1}),$
$\textrm{pr}(A=1\mid A^{*}=0,X)$ $=\text{logit}^{-1}(-1+0.5X_{1}-X_{2}),$
$\textrm{pr}(Z=1\mid A^{*}=1,X)$ $=\text{logit}^{-1}(-1-X_{1}+0.5X_{2}),$
$\textrm{pr}(Z=1\mid A^{*}=0,X)$ $=\text{logit}^{-1}(2+X_{1}).$
We consider two outcome regression models: one for a binary outcome and another for a continuous outcome.
i. (binary) $\textrm{pr}(Y=1\mid A^{*}=1,X)=\text{logit}^{-1}(-2+0.5X_{1}+X_{2}),$
$\textrm{pr}(Y=1\mid A^{*}=0,X)=\text{logit}^{-1}(1-2X_{1}+0.5X_{2})$
ii. (continuous) $Y=\sin(\pi X_{1})+(A^{*}-0.5)(X_{1}+X_{2})+\varepsilon,\quad\varepsilon\sim\mathcal{N}(0,1)$
In the M-step of Algorithm 1, we employ kernel regression, which has the form
$\begin{split}\widetilde{\textrm{pr}}(A^{*}=1\mid X=x)&=\frac{\sum^{n}_{i=1}K\left(\frac{x-x_{i}}{h}\right)p^{t}_{i}}{\sum^{n}_{i=1}K\left(\frac{x-x_{i}}{h}\right)}\\ \widetilde{\textrm{pr}}(A^{*}=1\mid Y=y,X=x)&=\frac{\sum_{i:Y_{i}=y}K\left(\frac{x-x_{i}}{h}\right)p^{t}_{i}}{\sum_{i:Y_{i}=y}K\left(\frac{x-x_{i}}{h}\right)}\text{ if $Y$ is binary}\\ \widetilde{\textrm{pr}}(A^{*}=1\mid Y=y,X=x)&=\frac{\sum_{i}K\left(\frac{\{(x-x_{i})^{2}+(y-y_{i})^{2}\}^{1/2}}{h}\right)p^{t}_{i}}{\sum_{i}K\left(\frac{\{(x-x_{i})^{2}+(y-y_{i})^{2}\}^{1/2}}{h}\right)}\text{ if $Y$ is continuous}\\ \widetilde{\textrm{pr}}(A^{*}=1\mid A=a,X=x)&=\frac{\sum_{i:A_{i}=a}K\left(\frac{x-x_{i}}{h}\right)p^{t}_{i}}{\sum_{i:A_{i}=a}K\left(\frac{x-x_{i}}{h}\right)}\\ \widetilde{\textrm{pr}}(A^{*}=1\mid Z=z,X=x)&=\frac{\sum_{i:Z_{i}=z}K\left(\frac{x-x_{i}}{h}\right)p^{t}_{i}}{\sum_{i:Z_{i}=z}K\left(\frac{x-x_{i}}{h}\right)}\end{split}$ (24)
where $K(\cdot)$ is a Gaussian kernel, and the bandwidth $h$ is calculated from the Silverman’s rule-of-thumb formula $h=1.06\text{sd}(X)n^{-1/(4+d)}$ (Silverman, 1986), with $d$ being the dimension of the regressor. Note that for $A$, we only need to estimate one of the probabilities $\textrm{pr}(A^{*}=1\mid A=a,X=x),a=0,1$, instead of both. This also applies to $\textrm{pr}(A^{*}=1\mid Z=z,X=x)$ for $z=0,1$ and $\textrm{pr}(A^{*}=1\mid Y=y,X=x)$ for $y=0,1$. In practice, one may elect to estimate the most prevalent surrogate subsample, i.e., we estimate $\textrm{pr}(A^{*}=1\mid A=1,X=x)$ if $n^{-1}\sum^{n}_{i=1}A_{i}\geq 0.5$, and $\textrm{pr}(A^{*}=1\mid A=0,X=x)$ if $n^{-1}\sum^{n}_{i=1}A_{i}<0.5$. We use 5-fold cross-fitting, and the number of Monte Carlo runs is 200. The threshold for convergence is set at 0.001.
For comparison, we also present results for an infeasible estimator which substitutes the true values of nuisance parameters into (20), and the naive estimator, which treats $A$ as $A^{*}$ and fits a logistic regression of $Y$ on $(A,X)$ for binary $Y$, and a kernel regression for continuous $Y$. In our setting, $A$ matches $A^{*}$ around 82% of the time. The results are displayed in Table 1 and 2. From the results, one can see that our proposed estimator shows robust performance, achieving coverage close to the nominal rate of 95%, while the naive estimator is severely biased.
@@
proposed infeasible naive proposed
$n=1000$ $\psi_{1}$ $0.01(0.17)$ $-0.19(0.14)$ $2.60(0.11)$ 94.5%
$\psi_{0}$ $-0.50(0.31)$ $-0.18(0.29)$ $-12.08(0.22)$ 97.5%
$n=2000$ $\psi_{1}$ $0.25(0.11)$ $0.04(0.10)$ $2.63(0.08)$ 96%
$\psi_{0}$ $-0.28(0.24)$ $-0.02(0.24)$ $-12.27(0.14)$ 94%
Table 1: Simulation results when $Y$ is binary. The “proposed” column is for our proposed estimator in (20). The coverage for the proposed estimator is given in the last column.
bias(s.e.)(×$\times 100$) coverage
proposed infeasible naive proposed
@@
Table 2: Simulation results when $Y$ is continuous. The “proposed” column is for our proposed estimator in (20). The coverage for the proposed estimator is given in the last column.
@@ 8 Real data example
Alzheimer’s Disease (AD) is a progressive neurodegenerative disorder that primarily affects the elderly, characterized by significant cognitive decline and memory impairment. One of the hallmark changes observed in AD is the atrophy of the hippocampus, a critical brain region involved in memory formation and spatial navigation. As the disease advances, neuronal damage occurs, resulting in brain atrophy that is especially pronounced in the hippocampus during the early stages of AD (Fox et al., 1996; Yu et al., 2022). We aim to analyze the impact of AD on hippocampal volume using data from the Alzheimer’s Disease Neuroimaging Initiative (ADNI). The data usage acknowledgment is included in §LABEL:supp-sec:adni_acknowledgment of the Supplementary Materials.
Diagnosing AD involves a comprehensive approach, combining cognitive assessments such as the Mini-Mental State Examination (MMSE) and neuroimaging techniques like MRI and PET scans to detect brain changes. However, AD diagnosis may not reflect the actual latent disease status (Hunter et al., 2015; Fischer et al., 2017). The diagnosis essentially serves as an approximation for the unobserved actual state of the disease. In practice, there are several other observable variables that can act as surrogates or proxies for the actual AD condition, such as behavioral scores based on standard questionnaires.
Our approach, as outlined in Section 5, requires selecting a surrogate and a proxy that are conditionally independent given a participant’s true but hidden AD status. Because AD diagnosis is essentially a composite of several measured factors including tests scores and biomarker measurements, using it as a surrogate would invalidate any such measurements as potential proxies given that they would fail to be independent. Instead, we use the recall and orientation subscores of the MMSE as indicators: one as a surrogate for underlying disease status and the other as a proxy. Specifically, the recall component evaluates the ability to remember three previously mentioned objects, while the orientation component assesses the awareness of the current time and place. We posit that the recall and orientation scores, which assess distinct cognitive functions, are independent given disease status and other covariates. This claim finds support in research conducted by Yew et al. (2013), which revealed distinct neural correlates of AD: performance on orientation scores was associated with the posterior hippocampus, whereas memory performance was more affected by the anterior hippocampal regions. The scores are converted into binary variables, where 0 indicates a perfect score in a subcategory, and 1 indicates any deviation. In our analysis, we also account for important observed confounders such as age, sex, and number of years of education (Vina and Lloret, 2010; Guerreiro and Bras, 2015; Zhang et al., 1990; Noble et al., 2012).
@@
We conduct separate analyses for the female and male groups, adjusting for age and years of education as potential confounders within each group. We consider the estimator proposed in equation (20) in §5. For the kernel function, we use the Gaussian kernel, with bandwidth selected using Silverman’s rule (Silverman, 1986). As in the simulation study, we use 5-fold cross-fitting. To address the influence of a particular random split on the estimates, we adopt the median adjustment method by repeating the cross-fitting 21 times (Chernozhukov et al., 2018).
Our results indicate that for females, AD leads to a decrease of 1,756 mm3 in hippocampal volume, with 95% confidence interval of $[1,370,2,142]$. For males, AD leads to a decrease of 1,406 mm3 in hippocampal volume, with 95% confidence interval of $[853,1,960]$ (Fox et al., 1996). These findings appear to be aligned with prior research indicating that females are affected by AD more severely than males, and AD-control brain weight differences were significantly greater for females (Filon et al., 2016). For comparison, we also performed linear regression for both groups, with hippocampal volume as the outcome, error-prone diagnosis as the main exposure, age, and education years as regressors. These analyses indicate an AD-related decrease of 1,162 mm3 for females (95% CI: $[976,1,347]$) and 971 mm3 for males (95% CI: $[793,1,151]$), which are of smaller magnitude than our estimates, suggesting the potential for effect attenuation, a common consequence of exposure measurement error. In addition to linear regression, we also implemented the more complex doubly robust Targeted Maximum Likelihood Estimator (TMLE) (Benkeser et al., 2017), which is essentially biased in this context and yields results similar to those of linear regression.

### v3

@@ 6 Extention to other functionals
    $\begin{split}\phi_{\mathrm{g}}^{F}(Y,A^{*},X;&\psi^{*})=\sum_{a^{*}=0,1}IF_{a^{*};\psi^{*}}=\sum_{a^{*}=0,1}\Bigl{\{}q_{0,a^{*}}(X)h_{0,a^{*}}(X)g_{1,a^{*}}(Y,X)\mathbb{I}(A^{*}=a^{*})\\ &+q_{0,a^{*}}(X)g_{2,a^{*}}(Y,X)\mathbb{I}(A^{*}=a^{*})+h_{0,a^{*}}(X)g_{3,a^{*}}(Y,X)+g_{4,a^{*}}(Y,X)\Bigr{\}},\end{split}$ (20)
    where $q_{0,a^{*}}(X)$ and $h_{0,a^{*}}(X)$ are unknown nuisance functions, $g_{j,a^{*}}$ are known functions of $(Y,X)$, and $IF_{a^{*};\psi^{*}},a^{*}=0,1$, belongs to the doubly robust class of influence functions studied in Robins et al., (2008). To simplify the exposition, we suppress the influence function dependence on $\psi^{*}$ throughout as $E[IF_{a^{*};\psi^{*}}]=0,a^{*}=0,1$. The next theorem characterizes for each influence function in the class above, a corresponding observed data influence function in the hidden treatment model. While we state the result for the class given above, the proof is more general as it considers an even larger class of doubly robust influence functions introduced in Ghassami et al., (2022), an important generalization which includes influence functions for causal functionals identified using the proximal causal inference framework (Miao et al.,, 2016, 2018; Tchetgen Tchetgen et al.,, 2024) that do not strictly fall within the above doubly robust class (Cui et al.,, 2024). The proof is deferred to the Supplementary Materials §LABEL:supp-sec:proof_extension.
@@
    Suppose that the influence functions for the parameter of interest $\psi^{*}$ in the full data model $\mathcal{M}^{*}$ are of the form in (20), indexed by functions $\mathrm{g}=(g_{1,a^{*}},g_{2,a^{*}},g_{3,a^{*}},g_{4,a^{*}},a^{*}=0,1)\in\mathcal{G}$ for a certain space of functions $\mathcal{G}$ in $L_{0}^{2}(X,Y)$ implied by the semiparametric model; then
@@
    is an influence function for $\psi^{*}$ in the hidden treatment model $\mathcal{M}$, where
    $\begin{split}M_{a^{*},A}=\{A-E(A\mid A^{*}=1-a^{*},X)\}\{E(A\mid A^{*}=a^{*},X)-E(A\mid A^{*}=1-a^{*},X)\}^{-1},\\ M_{a^{*},Z}=\{Z-E(Z\mid A^{*}=1-a^{*},X)\}\{E(Z\mid A^{*}=a^{*},X)-E(Z\mid A^{*}=1-a^{*},X)\}^{-1}.\end{split}$ (22)
@@
      Perform a one-step update in the direction of estimated EIF $\hat{\psi}^{eff}=\hat{\psi}+\mathbb{P}_{n}\hat{\phi}_{eff}(Y,A,Z,X),$ where $\hat{\phi}_{eff}(Y,A,Z,X)$ is an estimator of the EIF of ${\phi}_{eff}(Y,A,Z,X)=\Pi(\phi_{\mathrm{g}}\mid\Lambda)$ evaluated at $\hat{\psi}$ given in (LABEL:supp-eq:est_eif) in the Supplementary Materials §LABEL:supp-sec:closed_form_w.
@@
    Semi-linear structural regression model (continued). The full data influence function in (5) can be reformulated as $\phi^{F}_{sl}=\mathbb{I}(A^{*}=1)h(X)\{Y-\beta_{a^{*}}-\beta_{x}(X)\}\{1-\textrm{pr}(A^{*}=1\mid X)\}+\mathbb{I}(A^{*}=0)h(X)\{Y-\beta_{x}(X)\}\{-\textrm{pr}(A^{*}=1\mid X)\}.$ An influence function in the observed data model can be obtained by replacing $\mathbb{I}(A^{*}=a^{*}),a^{*}=0,1$ with $M_{a^{*},A}(A,X)M_{a^{*},Z}(Z,X)$ in (22), which gives
@@ 7 Simulation study
    In this section, we evaluate the performance of our proposed estimator for both binary and continuous outcomes. The covariates $X=(X_{1},X_{2})$ are generated from a 2-dimensional uniform $[0,1]^{2}$ distribution. We then generate $A^{*}$ conditional on $X$ from a Bernoulli distribution with $\textrm{pr}(A^{*}=1\mid X)=1/\{1+\exp(-X_{1})+\exp(-X_{2})\}$. Next, we generate $A$ and $Z$ from $\textrm{pr}(A=1\mid A^{*}=1,X)=\text{logit}^{-1}(1+2X_{1})$, $\textrm{pr}(A=1\mid A^{*}=0,X)=\text{logit}^{-1}(-1+0.5X_{1}-X_{2})$, $\textrm{pr}(Z=1\mid A^{*}=1,X)=\text{logit}^{-1}(-1-X_{1}+0.5X_{2})$, $\textrm{pr}(Z=1\mid A^{*}=0,X)=\text{logit}^{-1}(2+X_{1})$, where $\text{logit}(p)=\ln p(1-p)^{-1}$.
    We consider two outcome regression models:
    i. (binary outcome) $\textrm{pr}(Y=1\mid A^{*}=1,X)=\text{logit}^{-1}(-2+0.5X_{1}+X_{2}),$
    $\textrm{pr}(Y=1\mid A^{*}=0,X)=\text{logit}^{-1}(1-2X_{1}+0.5X_{2});$
    ii. (continuous outcome) $Y=\sin(\pi X_{1})+(A^{*}-0.5)(X_{1}+X_{2})+\varepsilon,\quad\varepsilon\sim\mathcal{N}(0,1).$
    The specific form for our kernel regression model in M-step is given in the Supplementary Materials §LABEL:supp-sec:simu_details. Note that for $A$, we only need to estimate one of the probabilities $\textrm{pr}(A^{*}=1\mid A=a,X=x),a=0,1$, instead of both. This also applies to $\textrm{pr}(A^{*}=1\mid Z=z,X=x)$ for $z=0,1$ and $\textrm{pr}(A^{*}=1\mid Y=y,X=x)$ for $y=0,1$. In practice, one may elect to estimate the most prevalent surrogate subsample, i.e., we estimate $\textrm{pr}(A^{*}=1\mid A=1,X=x)$ if $n^{-1}\sum^{n}_{i=1}A_{i}\geq 0.5$, and $\textrm{pr}(A^{*}=1\mid A=0,X=x)$ if $n^{-1}\sum^{n}_{i=1}A_{i}<0.5$. We use 5-fold cross-fitting, and the number of Monte Carlo runs is 200. The threshold for convergence is set at 0.001.
    For comparison, we also present results for an infeasible estimator which substitutes the true values of nuisance parameters into (19), and the naive estimator, which treats $A$ as $A^{*}$ and fits a logistic regression of $Y$ on $(A,X)$ for binary $Y$, and a kernel regression for continuous $Y$. In our setting, $A$ matches $A^{*}$ around 82% of the time. The results are displayed in Table 1. From the results, one can see that our proposed estimator shows robust performance, achieving coverage close to the nominal rate of 95%, while the naive estimator is severely biased.
@@
    $Y$ proposed infeasible naive proposed
    $n=1000$ $\psi_{1}$ $0.14(0.23)$ $-0.04(0.22)$ $8.49(0.15)$ 92.5%
    Binary $\psi_{0}$ $0.19(0.20)$ $0.13(0.19)$ $-4.15(0.15)$ 95%
    $n=2000$ $\psi_{1}$ $0.04(0.16)$ $-0.11(0.13)$ $8.51(0.10)$ 96%
    $\psi_{0}$ $-0.01(0.15)$ $-0.04(0.14)$ $-4.21(0.11)$ 95.5%
@@
    Table 1: Simulation results when $Y$ is binary or continuous. The “proposed” column is for our proposed estimator in (19). The “infeasible” column is for the estimator using true values of nuisance parameters. The “naive” column is for the estimator that treats $A$ as $A^{*}$. The coverage for the proposed estimator is given in the last column.
@@ 8 Real data example
    Alzheimer’s Disease (AD) is a progressive neurodegenerative disorder that primarily affects the elderly, characterized by significant cognitive decline and memory impairment. One of the hallmark changes observed in AD is the atrophy of the hippocampus, a critical brain region involved in memory formation and spatial navigation. As the disease advances, neuronal damage occurs, resulting in brain atrophy that is especially pronounced in the hippocampus during the early stages of AD (Fox et al.,, 1996; Yu et al.,, 2022). We aim to analyze the impact of AD on hippocampal volume using data from the Alzheimer’s Disease Neuroimaging Initiative (ADNI). The data usage acknowledgment is included in §LABEL:supp-sec:adni_acknowledgment of the supplement.
    Diagnosing AD involves a comprehensive approach, combining cognitive assessments such as the Mini-Mental State Examination (MMSE) and neuroimaging techniques like MRI and PET scans to detect brain changes. However, AD diagnosis may not reflect the actual latent disease status (Hunter et al.,, 2015). The diagnosis essentially serves as an approximation for the unobserved actual state of the disease.
    Our approach, as outlined in Section 5, requires selecting a surrogate and a proxy that are conditionally independent given a participant’s true but hidden AD status. Because AD diagnosis is essentially a composite of several measured factors including tests scores and biomarker measurements, using it as a surrogate would invalidate any such measurements as potential proxies given that they would fail to be independent. Instead, we use the recall and orientation subscores of the MMSE as indicators: one as a surrogate for underlying disease status and the other as a proxy. Specifically, the recall component evaluates the ability to remember three previously mentioned objects, while the orientation component assesses the awareness of the current time and place. We posit that the recall and orientation scores, which assess distinct cognitive functions, are independent given disease status and other covariates. This claim finds support in Yew et al., (2013), which revealed distinct neural correlates of AD: performance on orientation scores was associated with the posterior hippocampus, whereas memory performance was more affected by the anterior hippocampal regions. The scores are converted into binary variables, where 0 indicates a perfect score in a subcategory, and 1 indicates any deviation. In our analysis, we also account for important observed confounders such as age, sex, and number of years of education.
@@
    We conduct separate analyses for the female and male groups, adjusting for age and years of education as potential confounders within each group. We consider the estimator proposed in equation (19) in §5. As in the simulation study, we use 5-fold cross-fitting. To address the influence of a particular random split on the estimates, we adopt the median adjustment method by repeating the cross-fitting 21 times (Chernozhukov et al.,, 2018).
    Our results indicate that for females, AD leads to a decrease of 1,756 mm3 in hippocampal volume, with 95% confidence interval of $[1,370,2,142]$. For males, AD leads to a decrease of 1,406 mm3 in hippocampal volume, with 95% confidence interval of $[853,1,960]$ (Fox et al.,, 1996). These findings appear to be aligned with prior research indicating that females are affected by AD more severely than males, and AD-control brain weight differences were significantly greater for females (Filon et al.,, 2016). For comparison, we also performed linear regression for both groups, with hippocampal volume as the outcome, error-prone diagnosis as the main exposure, age, and education years as regressors. These analyses indicate an AD-related decrease of 1,162 mm3 for females (95% CI: $[976,1,347]$) and 971 mm3 for males (95% CI: $[793,1,151]$), which are of smaller magnitude than our estimates, suggesting the potential for effect attenuation, a common consequence of exposure measurement error. In addition to linear regression, we also implemented the more complex doubly robust Targeted Maximum Likelihood Estimator (Benkeser et al.,, 2017), which is essentially biased in this context and yields results similar to those of linear regression.